Controls Meeting 12-Jul-2012
- Minutes from previous meeting 28-Jun-2012
- JInventory - Josh, Hovanes, Elliott
- barcode scanner
- Controls hardware purchases - Josh, Hovanes
- EPICS - Hovanes, Vanik
- motion control, alarm system, goniometer, UConn boards, other power supplies, etc.
- Counting house installation - Hovanes, Mark, Elliott
- do we need to communicate with both UPS and RDC?
- Solenoid test - Mark, Josh, Scot, Elliott
- Fasttrack schedule
- Maintenance and other OSP's
- Goals, test plan, procedures, reviews
- FMEA, SOE, QD, overcurrent protection, ground fault detector, etc.
- PLC programming and PXI system - Josh, Dave, Yi
- Test stand - Hovanes, Josh
- Elogs and databases - Elliott
- Mantis task tracker
- Experiment Control Supervisor - Elliott
9:30 AM Thurs 12-Jul-2012 CC L207
9:30am Thurs 26-Jul-2012 CC L207
Present: Elliott W, Josh B, Mark S, Vanik K.
Hovanes tested a loaner barcode scanner and was not happy with it. We should be getting another. Josh entered a complete A-B chassis with slots and needs to make some minor modifications.
- Josh ordered the CompactLogix system for the BCAL RTD's (two systems plus 16 RTD modules).
- Still need to order modules for BCAL humidity sensors, start counter and CDC thermocouples.
- Josh ordered a quad-core computer to host the A-B license and other servers.
- Hovanes ordered some terminal servers, similar to the ones used in CLAS.
- Mark and Mark Baer are ordering parts for the Rack Room power distribution.
- Vanik recalls Hovanes ordered either some motors and/or a stage.
Vanik reported that the EPICS driver for the motion controller is complete, and implements all functions supported by the controller. In his last month here he will develop gui's for various devices that will use the controller. Also, Nerses is due to arrive soon and will work on HV and LV control.
- The racks are all in the control room.
- Mark needs to have some floor tiles cut.
- Otherwise the Rack Room is ready for power installation.
- Elliott will figure out if there's anything interesting to monitor in the RDC.
- Tim will import some of Tom's scheduling information into the FT schedule while Tom is in vacation.
- Giles will work on the Danfysik maintenance OSP.
- We will install one of our network switches in the controls racks.
- We have enough computers now (three Windows and one Linux box).
- We want the computers to be located on the same level as the controls racks. Elliott will inform Tom.
- Elliott sent out a preliminary version of goals for the test. This will lead to a test plan, followed by detailed procedure documents.
- We are still on track for a review in October.
- George continues to work on the FMEA and vacuum system.
- Moving the Danfysik has been delayed until the fire pipes are installed (next week).
PLC and PXI systems
- Josh is well along reprogramming the PLC.
- Dave has been busy with other things but should be ready in time.
- Yi has not started yet but should have no problem finishing in time.
- Josh received the second SOE module.
- Mark needs to set up a two-SOE system for Josh to test.
Elog and Databases
- Theo put out a prototype for the new ELog based on Drupal.
- New system supposed to address requirements from all four halls.
- With luck we'll be able to use it for the solenoid test.
- Mark and Josh have put many manuals into the Midas system.